OpenStudy (gabylovesyou):

Find the height of a square pyramid that has a volume of 12 cubic feet and a base length of 3 feet. 1.5 feet 2 feet 3 feet 4 feet

OpenStudy (gabylovesyou):

@jim_thompson5910

ganeshie8 (ganeshie8):

use volume of pyramid formula

OpenStudy (gabylovesyou):

1/2bh ?

ganeshie8 (ganeshie8):

\[V = \dfrac{1}{3}Bh\] \(B \) = Area of base \(h\) = height of pyramid

OpenStudy (gabylovesyou):

12 = 1/3 * 3 * h

ganeshie8 (ganeshie8):

its a square pyramid - that means the base is a square

ganeshie8 (ganeshie8):

and you're given that the length of square is 3 units so whats the Area of square whose side is 3 units ?

OpenStudy (gabylovesyou):

3 * 3 = 9 ?

ganeshie8 (ganeshie8):

Yes ! so \(B\) = 9

ganeshie8 (ganeshie8):

plug that in the formula ans solve h

OpenStudy (gabylovesyou):

4 ?

ganeshie8 (ganeshie8):

\[V = \dfrac{1}{3}Bh\] \[12 = \dfrac{1}{3}(9)h\] \[12 = 3h\]

ganeshie8 (ganeshie8):

yes, h = 4 !

OpenStudy (gabylovesyou):

thank you!

ganeshie8 (ganeshie8):

yw